X^2-2X-8y+17=0 Find the vertex, focus, equations of the directrix, and the axis of symmetry. Please write each step out. I can complete the square and then put it in standard form y=ax^2. My question is identifying all the above from the standard form.
Answer by stanbon(75887)

X^2-2X-8y+17=0 Find the vertex, focus, equations of the directrix, and the axis of symmetry. Please write each step out. I can complete the square and then put it in standard form y=ax^2. My question is identifying all the above from the standard form.
----------
8y = x^2-2x+17
-----
x^2 - 2x = 8y-17
x^2-2x+1 = 8y-16
(x-1)^2 = 2(y-2)
-----
Vertex:: (1,2)
Axis of symmetry:: x = 1
4p = 2
p = 1/2
------
focus:: (1,2+p) or (1,5/2)
directrix: y = 2-p = 3/2
-------
Cheers,
Stan H.